OpenStudy (anonymous):

let f(x)=1-2x^2 and g(x) = x+1 compute the following (f o g)(-1)

OpenStudy (anonymous):

the answr is 1

OpenStudy (anonymous):

I got the correct answer I just need to make sure my steps are correct.

OpenStudy (anonymous):

fg(x) is also (f o g)(x) but I'll just write it my way. \[fg(x)=-2x^{2}-4x-1\]\[fg(-1)=-2(-1)^{2}-4(-1)-1\]\[fg(-1)=1\]

OpenStudy (anonymous):

where are you getting the 4

OpenStudy (anonymous):

f(g(x)) = 1-2(x+1)^2 = 1-2*0^2 =1

OpenStudy (anonymous):

the steps are above

OpenStudy (anonymous):

ok I got it thanks.